OpenStudy (anonymous):

What is the equation of the parabola whose vertex lies on the coordinates (12, 7

OpenStudy (anonymous):

Seems like you are missing some information. Is that all that was given?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You could use vertex form: f(x) = a(x - h)^2 +k where (h, k) is the vertex. I guess just make a = 1.

OpenStudy (anonymous):

y= 1(12-14)^2 +7?

OpenStudy (anonymous):

y=(x-12)^2 +7